OpenStudy (anonymous):

Solve the system. 5x+8y=41 x-y=3 (x,y)=

OpenStudy (katrinakaif):

Lets use substitution here ;)

OpenStudy (katrinakaif):

In your second equation "x - y = 3" can be re-written as x = 3 + y Correct?

OpenStudy (katrinakaif):

Now lets substitute the concluded equation "x = 3 + y" into the equation "5x + 8y = 41"

OpenStudy (katrinakaif):

That'd give us: 5 (3 + y) + 8y = 41

OpenStudy (katrinakaif):

Now simply simplify. Distribute 5 to (3 + y) giving us : 15 + 5y + 8y = 41 Then collect like terms.. 15 + 13y = 41 Bring 15 to the other side by substracting.. 13y = 41 - 15 13y = 26 Divide by 13 on both sides y = 2

OpenStudy (anonymous):

if y=2 what does x=

OpenStudy (katrinakaif):

Now substitute y = 2 into the original equation "5x + 8y = 41" giving us 5x + 8(2) = 41 Now simplify this, and you will get your "x" Simply arrange them into the ordered pair in the designated places, as (x, y)

OpenStudy (katrinakaif):

Sorry sweetie, I wont be completing the work for what "x" equals. Give it a try, and if you need help, I shall guide you through it. =)

OpenStudy (anonymous):

x=5

OpenStudy (katrinakaif):

Correct! Great work ;)

OpenStudy (katrinakaif):

Our last step is listing this as an ordered pair..

OpenStudy (katrinakaif):

and that would conclude as our final answer =)

OpenStudy (anonymous):

ok